A ball is kicked upward with an initial velocity of 56 feet per second. The ball's height h (in feet) can be expressed as a function of time t (in seconds) by the equation h = -16t2 + 56t. How much time does the ball take to reach its highest point?
Answer by nerdybill(7384)

A ball is kicked upward with an initial velocity of 56 feet per second. The ball's height h (in feet) can be expressed as a function of time t (in seconds) by the equation h = -16t2 + 56t. How much time does the ball take to reach its highest point?
.
Max is at the vertex.
t value at the vertex can be found by:
t = -b/(2a)
.
substitute in the values of a,b from:
h = -16t2 + 56t
.
t = -b/(2a)
t = -56/(2(-16))
t = -56/(-32)
t = 1.75 seconds